OpenStudy (anonymous):

If x is a whole number and 437 = (21 + x)(21 - x), then x = __________

OpenStudy (anonymous):

@calculusxy @ganeshie8 @Michele_Laino

OpenStudy (anonymous):

@Vocaloid

imqwerty (imqwerty):

1st we open the brackets \[437=21^2-x^2\]\[437=441-x^2\]\[x^2=4\]\[x=\pm 2\]but since x is a whole number x= +2